# Three Scores and 15 Years (1948-2023) of Rao’s Score Test: A Brief History

### v1

@@ Abstract
Rao, (1948) introduced the score test statistic as an alternative to the likelihood ratio and Wald test statistics. In spite of the optimality properties of the score statistic shown in Rao and Poti, (1946), the Rao score (RS) test remained unnoticed for almost 20 years. Today, the RS test is part of the “Holy Trinity” of hypothesis testing and has found its place in the Statistics and Econometrics textbooks and related software. Reviewing the history of the RS test we note that remarkable test statistics proposed in the literature earlier or around the time of Rao, (1948) mostly from intuition, such as Pearson, (1900) goodness-fit-test, Moran, (1948) I test for spatial dependence and Durbin and Watson, (1950) test for serial correlation, can be given RS test statistic interpretation. At the same time, recent developments in the robust hypothesis testing under certain forms of misspecification, make the RS test an active area of research in Statistics and Econometrics. From our brief account of the history the RS test we conclude that its impact in science goes far beyond its calendar starting point with promising future research activities for many years to come.
Key words: Hypothesis Testing; Rao’s Score; Applications to Econometrics and Statistics; Sequential Testing; Robust Tests.
@@
C. R. Rao’s work was always inspired by some practical problems. In 1946, he was deputed from the Indian Statistical Institute (ISI), Calcutta, to work on an anthropometric project in the Museum of Anthropology and Ethnology at the Cambridge University, U.K. While at Cambridge, Rao took the opportunity to contact R. A. Fisher, then the Belfour Professor of Genetics, and registered for a Ph.D. degree in Statistics under Fisher’s guidance. As recollected in Rao, (2001), Fisher agreed under the condition that Rao spends time in the Genetics Laboratory where Fisher was breeding mice to map their chromosomes. Rao started by mating mice of different genotypes to collect the necessary data and additionally, he was trying to develop the appropriate statistical methodology to analyze the experimental data. The problem was estimation of linkage parameters (recombination probabilities in the various segments of the chromosomes) using data sets from different experiments, designed in such a way that each data set had information on the same parameters. It was thus necessary to test whether the parameters in different experiments are the same or not.
@@
Rao wrote and published two papers based on this work. The first paper dealt with the general problem of testing simple and composite hypotheses concerning a vector parameter. The test was based on the scores, derivatives of the log-likelihood function with respect to the individual parameters. The paper was published in the Proceedings of the Cambridge Philosophical Society, [Rao, (1948)] where he termed the test principle as a score test. In this paper, we will refer to it as the Rao score (RS) test. The other paper, contained the detail steps for analyzing the data involving the segregation of several factors in mating of different genotypes. And it used the RS test for the meta-analysis of testing the equality of parameters coming from different experimental data sets. That paper was published in Fisher’s new journal Heredity, Rao, 1950a . For more, see Rao, (2001).
@@
The rest of the paper is organized as follows. In Section 2, we start with the first principle of testing, namely the Neyman-Pearson Lemma and derive the simplest version of RS test and then discuss it in its full generality. There we also provide RS test interpretation to some of the classic tests in Econometrics and Statistics, such as the quintessential Pearson, (1900) goodness-fit-test, which was suggested mostly by pure intuition, but its theoretical foundation can be buttressed by RS test principle. In Sections 3 and 4, we list a (somewhat incomplete) catalogue of RS tests in Econometrics and Statistics. In Section 5, we outline some of the possible ways an assumed probability model can be misspecified, and discuss how the various RS tests can be robustified to make them valid under misspecification. We close the paper in Section 6 (Epilogue) with some concluding remarks. At the outset let us mention that while compiling the 75 years (from 1948 to 2023) history of the RS test, we have included here some of our own past historical accounts and cited accordingly. Our aim was to have a comprehensive review, as far as possible, at one place, like a one-stop-shopping for the RS test history.
@@
We start by introducing some notation and concepts. Suppose we have $n$ independent observations $y_{1},y_{2},...,y_{n}$ on a random variable $Y$ with density function $f(y;\theta)$, where $\theta$ is a $p$ x $1$ parameter vector with $\theta\in\Theta\in\mathbb{R}^{p}$. It is assumed that $f(y;\theta)$ satisfies the regularity conditions stated in Rao, (1973, p.364) and Serfling, (1980, p.144). The likelihood function is given by
@@
where y $=(y_{1},y_{2},...,y_{n})$ denotes the sample. Suppose we want to test $H_{0}:\theta=\theta_{0}$ against $H_{1}:\theta\neq\theta_{0}$ based on the sample y.
The foundation of the theory of hypothesis testing was laid by Neyman and Pearson, (1933) fundamental lemma. This lemma provides a way to find the most powerful (MP) and uniformly most powerful (UMP) tests. According to the Neyman-Pearson (N-P) Lemma, the MP critical region having size $\alpha$, is given by
@@
[…]
@@
Using the generalized N-P Lemma given in Neyman and Pearson, (1936), it is easy to see that the locally most powerful (LMP) test for $H_{0}:\theta=\theta_{0}$ vs $H_{1}:\theta>\theta_{0}$, is given by
$\frac{\partial}{\partial\theta}\mathbf{L}(\theta_{0})>\kappa\mathbf{L(\theta_{0})}$
@@
$\frac{\partial}{\partial\theta}\ln(\mathbf{L}(\theta_{0}))=\frac{\partial}{\partial\theta}l(\theta_{0})>\kappa,$ (6)
where $l(\theta)$ denotes the log-likelihood function and $\kappa$ is a constant such that the size of the test is $\alpha$. The quantity $S(\theta)={\partial l(\theta)}/{\partial\theta}$ is called the Fisher-Rao score function. The above result in (6) was first discussed in Rao and Poti, (1946), who stated that a LMP test for $H_{0}:\theta=\theta_{0}$ is given by
$\kappa_{1}S(\theta_{0})>\kappa_{2}$ (7)
@@
2.1.From Rao and Poti, (1946) to Rao, (1948): Test for the Multiparamter Case
@@
which under $H_{0}$ will be distributed as $\chi^{2}_{1}$ as in Rao and Poti, (1946). Note that our $H_{0}:\theta=\theta_{0}$ for $p\geq 2$ can be expressed as $H_{0}\equiv\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{0\delta}$, i.e., the multi parameter testing problem can be decomposed into a series of single-parameter problems. To obtain a linear function like (9), Rao, (1948) maximized (11) with respect to $\delta$. Using Cauchy-Schwarz inequality
@@
Roy (1953) used Rao’s maximization technique (13) to develop his union-intersection (UI) method of testing. Let $H_{0}\equiv\underset{j\in J}{\bigcap}H_{0j}$, where J is an index set. Roy’s UI method gives the rejection region for $H_{0}$ as the union of rejection regions for all $H_{0j}$, $j\in J$. Consider testing $H_{0\delta}:$ $\delta^{\prime}\theta=\delta^{\prime}\theta_{0}$ against $H_{1\delta}:$ $\delta^{\prime}\theta\neq\delta^{\prime}\theta_{0}$, $\delta\in\mathbb{R}^{p}$. Let $H_{0}=\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{0\delta}$ and $H_{1}\equiv\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{1\delta}$. If $T_{\delta}$ is the likelihood ratio (LR) statistic for testing $H_{0\delta}$ against $H_{1\delta}$, then
@@
where $\widetilde{\theta}$ is the restricted maximum likelihood estimator (MLE) of $\theta$, i.e., $h(\tilde{\theta})=c$. Asymptotically, under $H_{0}$, the RS test statistic is distributed as $\chi^{2}_{r}$. Therefore, we observe two optimality principles behind the RS test; first, in terms of LMP test as given in (6), and second, in deriving the “optimal” direction for the multiparamter case, as in (13).
@@
Their suggestion did not come from any search procedure satisfying an optimality criterion, it was purely based on intuitive grounds, as Neyman (1980, p.6) stated, “The intuitive background of the likelihood ratio test was simply as follows: if among the contemplated admissible hypotheses there are some that ascribe to the facts observed probabilities much larger than that ascribed by the hypothesis tested, then it appears ‘reasonable’ to reject the null hypothesis.”
The three statistics RS, W, and LR, given respectively in (15), (16), and (17) are referred to as the “Holy Trinity.” These tests can be viewed as three different distance measures between $H_{0}$ and $H_{1}$. When $H_{0}$ is true, we should expect the restricted and unrestricted MLEs of $\theta$, $\widetilde{\theta}$, and $\hat{\theta}$ to be close, and likewise the log-likelihood functions $l(\widetilde{\theta})$ and $l(\hat{\theta})$, respectively. The LR statistic in (17) measures the distance through the log-likelihood function and is based on the difference $l(\hat{\theta})-l(\widetilde{\theta})$. To see the intuition behind the RS test, note that $S(\hat{\theta})=0$ by construction, and thus we should expect $S(\widetilde{\theta})$ to be cloes to zero if $H_{0}$ is true. Therefore, the basis of the RS test is $S(\widetilde{\theta})-S(\hat{\theta})=S(\widetilde{\theta})$, distance between $\widetilde{\theta}$ and $\hat{\theta}$ measured through the function $S(\theta)$. Finally to test $H_{0}:h(\theta)=c$, W considers the distance directly in terms of $h(\theta)$, and is based on $\left[h(\hat{\theta})-c\right]-\left[h(\widetilde{\theta})-c\right]=h(\hat{\theta})-c$, where $h(\widetilde{\theta})=c$ by construction, as we see in expression (16). It is interesting to note the duality between the Wald and the RS tests, respectively based on $h(\hat{\theta})$ and $S(\widetilde{\theta})$. Therefore the RS test statistic is closer to W than LR. Therefore, it makes sense that Rao (1948, p.53) mentioned his test as an alternative to W.
@@
From above, we can note the following features of the three tests. First, since the tests are based on three different areas in general, they will yield conflicting inference if the same critical value is used [see Berndt and Savin, (1977)]. Second, the RS tests depends only on $S(\theta)$ and the slope of $S(\theta)$ at $\theta_{0}$. We can draw many curved lines through F with the same slope at F, and the dotted line $A^{\prime}FD^{\prime}$ is an example. This implies that there may be other $S(\theta)$ functions, i.e., other likelihood function representing different alternative hypothesis, with the same slope at $\theta_{0}$, giving rise to the same RS test statistic. In the literature, this property is known as invariance property of the RS test principle [see Godfrey, (1988, p.70)]. Finally, for both the RS and W tests, the variances can be calculated in a number of ways which are asymptotically equivalent. This can lead to different versions of the test statistics. It is not clear which versions will give better results in a finite samples.
@@
RS test was well ahead of its time. It went unnoticed for very many years. It is fair to say that econometricians can claim major credit in recognizing its importance and applying the RS test in several useful contexts and coming up with closed form, neat test statistics. More recently, statisticians are catching up with innovative applications. To obtain a quantitative perception of the influence of Rao, (1948), we plot the yearly citations for the last 75 years in Figure 2. The corresponding cumulative citations are depicted in Figure 3. First thing to note is that the total number of citations in the last 75 years is only 980, apparently a very low number for such a seminal paper. Of course, we need to take into consideration of the fact that there are many papers, especially in the Statistics literature, that use the score test without making any reference to Rao, (1948). Second, there are only a handful citations during the first thirty years, i.e., until around 1978. That was the time econometricians recognized the usefulness of the Rao test principle, and used it in developing several model specification tests. There was another surge in its use after another 30 years, i.e., around 2008, in both the Statistics and Econometrics literature. Finally, from both Figures 2 and 3, it is clear that overall, the number of citations is still going up at an increasing rate, indicating continuing influence of Rao, (1948), as far as the citation numbers go.
@@
$E\left[\sqrt{b_{1}}\right]=0,\quad Var\left[\sqrt{nb_{1}}\right]=6$ (40)
@@
$P_{N}^{\prime}(\theta_{0})\geq\mu\,P_{N}(\theta_{0})$ (48)
@@
$P_{n}^{\prime}(\theta_{0})\geq A(N)\,P_{n}(\theta_{0})$ (49)
@@
Sequential testing procedures that perform interim analyses during the evolution of the experiment, with the goal of obtaining the result earlier than the termination time suggested by the fixed sample analysis due to time or monetary cost considerations or ethical reasons, are easier validated with the use of score-based test statistics rather than the analogues of LR statistics. We may refer to chapters 9-11 of Sen, (1981) for the role of score processes in sequential nonparametrics and where it is mentioned (p.339) “it is comparatively simpler to verify these regularity conditions [i.e., for the score] than those for the likelihood function.”
@@
To diagnose the potential ability of candidate judges and to decide on the selection of a taste panel, each candidate judge is required to perform a prespecified number of sample comparisons. The sufficient number of these sample comparisons before reaching to a decision on who to select for the panel is always a concern.
@@
Bilias, (2000) offered an application of a repeated significance test in a retrospective analysis of the Pennsylvania ‘Reemployment Bonus’ controlled experiments conducted by the US Department of Labor. Their main purpose was to determine whether the offer of a bonus amount to the unemployment insurance (UI) claimants, provided that they find a job with some required permanence within a given period of time, can act as an incentive for more intensive job-seeking with subsequent reduction of the unemployment spells. The response of primary interest is the length of insured unemployment spell and it is assumed that it follows the proportional hazard regression model. The statistic for measuring the effect of the various bonus packages on the duration of insured unemployment relative to the existing scheme is the partial likelihood score statistic. In carrying out sequential analysis, the score statistic is evaluated repeatedly, at different points in chronological time, each time with the available data. The retrospective sequential analysis concluded that the experiment could be concluded earlier than the fixed sample analysis with gains in time and monetary savings.
@@
While discussing the problem in statistical hypothesis testing, Haavelmo, (1944, pp.65-66) stated, “Whatever be the principles by which we choose a “best” critical region of size $\alpha$, the essential thing is that a test is always devloped with respect to a given fixed set of possible alternatives $\Omega^{0}$.” Haavelmo called $\Omega^{0}$, the priori admissible hypotheses and according to him, a test is not robust if we shift our attention to another admissible set $\Omega^{\prime}$ (that may be obtained by extending $\Omega^{0}$ to include new/different alternatives), for which the proposed test has poor size and power properties.
Very often it is difficult to interpret the results of a test applied to a misspecified model. For instance, while testing the significance of some of the regression coefficients in the linear regression models, the results are not easily interpretable when a nonlinear model is the appropriate one [see, White, (1980), Bera and Byron, (1983) and Byron and Bera, (1983)]. In the Statistics and Econometrics literature, most emphasis has been put on the minimization of type-I and type-II error probabilities. There are, however, only a few works that seriously consider the consequences and suggest remedies of misspecifying the priori admissible hypothesis – which can be called the type-III error.
Note that the model under our priori admissible hypothesis could be misspecified in a variety of ways. Here we consider only two kinds: distributional and parametric. In the former case, the assumed probability density function differs from the true data generating process (DGP). Kent, (1982) and White, (1982) analyzed this case and suggested modified version of the RS test that involves adjustment of the variance of the score function. In the parametric misspecification case, the dimension of the assumed parameter space does not match with the true one. Bera and Yoon, (1993) developed a modified RS test that is valid under the local parametric misspecification.
@@
Due to this divergence between $J$ and $K$, and noting that we defined the information matrix $\mathcal{I}(\theta)$ in (10) by taking expectation under $f(y;\theta)$ instead of under the DGP $g(y)$, the standard RS test in (15) is not valid. Kent, (1982) and White, (1982) suggested the following robust form of the RS test statistic for testing the $H_{0}:h(\theta)=c$:
@@
where $H(\theta)=\partial h(\theta)/\partial\theta$, $B(\theta)=K(\theta)^{-1}J(\theta)K(\theta)^{-1}$ and $\widetilde{\theta}$ denotes quasi-MLE (QMLE) of $\theta$, and the notation $RS^{*}(D)$ is used to signify robust RS test statistic under distributional misspecification.
Under $H_{0}:h(\theta)=c$, $RS^{*}(D)$ is asymptotcally distributed as $\chi^{2}_{r}$ even under distributional misspecification, that is, when the assumed density $f(y;\theta)$ does not coincide with the true DGP $g(y)$. This approach of finding the asymptotically correct formula for variance has its origin in Koopmans et al., (1950, pp.148-150); for more on this see Bera et al., (2021). Expression (56) can be simplified if the parameter vector $\theta$ $(p\times 1)$ can be decomposed as $\theta=(\gamma^{\prime},\psi^{\prime})^{\prime}$ where $\gamma$ and $\psi$ have dimensions $m$ and $r$, respectively, $m+r=p$, and we test $H_{0}:\psi=\psi_{*}$ (say). Let us also partition the score function $S(\theta)$ and $J(\theta)$ [similarly $K(\theta)$] as

### v2

# Three Score and 15 Years (1948-2023) of Rao’s Score Test: A Brief History

@@ Abstract
Rao, (1948) introduced the score test statistic as an alternative to the likelihood ratio and Wald test statistics. In spite of the optimality properties of the score statistic shown in Rao and Poti, (1946), the Rao score (RS) test remained unnoticed for almost 20 years. Today, the RS test is part of the “Holy Trinity” of hypothesis testing and has found its place in Statistics and Econometrics textbooks and related software. Reviewing the history of the RS test we note that remarkable test statistics proposed in the literature earlier or around the time of Rao, (1948) mostly from intuition, such as Pearson, (1900) goodness-fit-test, Moran, (1948) I test for spatial dependence and Durbin and Watson, (1950) test for serial correlation, can be given RS test statistic interpretation. At the same time, recent developments in the robust hypothesis testing under certain forms of misspecification, make the RS test an active area of research in Statistics and Econometrics. From our brief account of the history of the RS test we conclude that its impact in science goes far beyond its calendar starting point with promising future research activities for many years to come.
Key words: Applications to Econometrics and Statistics; Hypothesis Testing; Rao’s Score; Robust Tests; Sequential Testing.
@@
C. R. Rao’s work was always inspired by some practical problems. In 1946, he was deputed from the Indian Statistical Institute (ISI), Calcutta, to work on an anthropometric project in the Museum of Anthropology and Ethnology at the Cambridge University, U.K. While at Cambridge, Rao took the opportunity to contact R. A. Fisher, then the Belfour Professor of Genetics, and registered for a Ph.D. degree in Statistics under Fisher’s guidance. As recollected in Rao, (2001), Fisher agreed under the condition that Rao spends time in the Genetics Laboratory where Fisher was breeding mice to map their chromosomes. Rao started by mating mice of different genotypes to collect the necessary data and additionally, he was trying to develop appropriate statistical methodology to analyze the experimental data. The problem was estimation of linkage parameters (recombination probabilities in the various segments of the chromosomes) using data sets from different experiments, designed in such a way that each data set had information on the same parameters. It was thus necessary to test whether the parameters in different experiments are the same or not.
@@
Rao wrote and published two papers based on this work. The first paper, Rao, (1948), deals with the general problem of testing simple and composite hypotheses concerning a vector parameter. The test was based on the scores, derivatives of the log-likelihood function with respect to the individual parameters. The paper was published in the Proceedings of the Cambridge Philosophical Society, where he termed the test principle as a score test. In this paper, we will refer to it as the Rao score (RS) test. The other paper, Rao, 1950a , contains the detail steps for analyzing the data involving the segregation of several factors in mating of different genotypes. And it used the RS test for the meta-analysis of testing the equality of parameters coming from different experimental data sets. That paper was published in Fisher’s new journal Heredity. For more, see Rao, (2001).
@@
The rest of the paper is organized as follows. In Section 2, we start with the first principle of testing, namely the Neyman-Pearson Lemma and derive the simplest version of RS test and then discuss it in its full generality. There, we also provide RS test interpretation to some of the classic tests in Econometrics and Statistics, such as the quintessential Pearson, (1900) goodness-fit-test, which was suggested mostly by pure intuition, but its theoretical foundation can be buttressed by RS test principle. In Sections 3 and 4, we list a (somewhat incomplete) catalogue of RS tests in Econometrics and Statistics. In Section 5, we outline some of the possible ways an assumed probability model can be misspecified, and discuss how the various RS tests can be robustified to make them valid under misspecification. We close the paper in Section 6 (Epilogue) with some concluding remarks. At the outset let us mention that while compiling the 75 years (from 1948 to 2023) history of the RS test, we have included here some of our own past historical accounts and cited accordingly. Our aim is to have a comprehensive review as far as possible at one place, like a one-stop-shopping for the RS test history.
@@
We start by introducing some notation and concepts. Suppose we have $n$ independent observations $y_{1},y_{2},...,y_{n}$ on a random variable $Y$ with density function $f(y;\theta)$, where $\theta$ is a $p\times 1$ parameter vector with $\theta\in\Theta\subseteq\mathbb{R}^{p}$. It is assumed that $f(y;\theta)$ satisfies the regularity conditions stated in Rao, (1973, p.364) and Serfling, (1980, p.144). The likelihood function is given by
@@
where y $=(y_{1},y_{2},\ldots,y_{n})$ denotes the sample. Suppose we want to test $H_{0}:\theta=\theta_{0}$ against $H_{1}:\theta\neq\theta_{0}$ based on the sample y.
The foundation of the theory of hypothesis testing was laid by Neyman and Pearson, (1933) fundamental lemma. This lemma provides a way to find the most powerful (MP) and uniformly most powerful (UMP) tests. According to the Neyman-Pearson (N-P) Lemma, the MP critical region for testing $H_{0}:\theta=\theta_{0}$ versus $H_{1}:\theta=\theta_{1}$ having size $\alpha$, is given by
@@
where $\kappa$ is such that $\Pr[{\omega(\textbf{y})|H_{0}}]=\alpha$.
@@
Using the generalized N-P Lemma given in Neyman and Pearson, (1936), it is easy to see that the locally most powerful (LMP) critical region for $H_{0}:\theta=\theta_{0}$ versus $H_{1}:\theta>\theta_{0}$, is given by
$\omega(\textbf{y})=\left\{\textbf{y}\;\Big{|}\;\frac{\partial}{\partial\theta}\mathbf{L}(\theta_{0})>\kappa\mathbf{L}(\theta_{0})\right\}$
@@
$\omega(\textbf{y})=\left\{\textbf{y}\;\Big{|}\;\frac{\partial}{\partial\theta}\ln(\mathbf{L}(\theta_{0}))=\frac{\partial}{\partial\theta}l(\theta_{0})>\kappa\right\},$ (6)
where $l(\theta)$ denotes the log-likelihood function and $\kappa$ is a constant such that the size of the test is $\alpha$. The quantity $S(\theta)={\partial l(\theta)}/{\partial\theta}$ is called the Fisher-Rao score function. The above result in (6) was first discussed in Rao and Poti, (1946), who stated that a LMP critical region for $H_{0}:\theta=\theta_{0}$ is given by
$\omega(\textbf{y})=\{\textbf{y}\;|\;\kappa_{1}S(\theta_{0})>\kappa_{2}\},$ (7)
@@
2.1.From Rao and Poti, (1946) to Rao, (1948): Test for the Multiparameter Case
@@
which under $H_{0}$ will be distributed as $\chi^{2}_{1}$ as in Rao and Poti, (1946). Note that our $H_{0}:\theta=\theta_{0}$ for $p\geq 2$ can be expressed as $H_{0}\equiv\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{0\delta}$, i.e., the multiparameter testing problem can be decomposed into a series of single-parameter problems. To obtain a linear function like (9), Rao, (1948) maximized (11) with respect to $\delta$. Using the Cauchy-Schwarz inequality
@@
Roy, (1953) used Rao’s maximization technique (13) to develop his union-intersection (UI) method of testing. Let $H_{0}\equiv\underset{j\in J}{\bigcap}H_{0j}$, where $J$ is an index set. Roy’s UI method gives the rejection region for $H_{0}$ as the union of rejection regions for all $H_{0j}$, $j\in J$. Consider testing $H_{0\delta}:$ $\delta^{\prime}\theta=\delta^{\prime}\theta_{0}$ against $H_{1\delta}:$ $\delta^{\prime}\theta\neq\delta^{\prime}\theta_{0}$, $\delta\in\mathbb{R}^{p}$. Let $H_{0}=\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{0\delta}$ and $H_{1}\equiv\underset{\delta\in\mathbb{R}^{p}}{\bigcap}H_{1\delta}$. If $T_{\delta}$ is the likelihood ratio (LR) statistic for testing $H_{0\delta}$ against $H_{1\delta}$, then
@@
where $\widetilde{\theta}$ is the restricted maximum likelihood estimator (MLE) of $\theta$, i.e., $h(\tilde{\theta})=c$. Asymptotically, under $H_{0}$, the RS test statistic is distributed as $\chi^{2}_{r}$. Therefore, we observe two optimality principles behind the RS test; first, in terms of LMP test as given in (6), and second, in deriving the “optimal” direction for the multiparameter case, as in (13).
@@
Their suggestion did not come from any search procedure satisfying an optimality criterion. It was purely based on intuitive grounds; as Neyman, (1980, p.6) stated, “The intuitive background of the likelihood ratio test was simply as follows: if among the contemplated admissible hypotheses there are some that ascribe to the facts observed probabilities much larger than that ascribed by the hypothesis tested, then it appears ‘reasonable’ to reject the null hypothesis.”
The three statistics RS, W, and LR, given respectively in (15), (16), and (17) are referred to as the “Holy Trinity.” These tests can be viewed as three different distance measures between $H_{0}$ and $H_{1}$. When $H_{0}$ is true, we should expect the restricted and unrestricted MLEs of $\theta$, namely $\widetilde{\theta}$ and $\hat{\theta}$ to be close, and likewise the log-likelihood functions $l(\widetilde{\theta})$ and $l(\hat{\theta})$, respectively. The LR statistic in (17) measures the distance through the log-likelihood function and is based on the difference $l(\hat{\theta})-l(\widetilde{\theta})$. To see the intuition behind the RS test, note that $S(\hat{\theta})=0$ by construction, and thus we should expect $S(\widetilde{\theta})$ to be cloes to zero if $H_{0}$ is true. Therefore, the basis of the RS test is $S(\widetilde{\theta})-S(\hat{\theta})=S(\widetilde{\theta})$, distance between $\widetilde{\theta}$ and $\hat{\theta}$ measured through the function $S(\theta)$. Finally to test $H_{0}:h(\theta)=c$, W considers the distance directly in terms of $h(\theta)$, and is based on $\left[h(\hat{\theta})-c\right]-\left[h(\widetilde{\theta})-c\right]=h(\hat{\theta})-c$, where $h(\widetilde{\theta})=c$ by construction, as we see in expression (16). It is interesting to note the similarity between the Wald and the RS tests based on $h(\hat{\theta})$ and $S(\widetilde{\theta})$, respectively. Therefore the RS test statistic is closer to W than LR. Therefore, it makes sense that Rao, (1948, p.53) mentioned his test as an alternative to W.
@@
From above, we can note the following features of the three tests. First, since the tests are based on three different areas in general, they will yield conflicting inference if the same critical value is used [see Berndt and Savin, (1977)]. Second, the RS tests depends only on $S(\theta)$ and the slope of $S(\theta)$ at $\theta_{0}$. We can draw many curved lines through F with the same slope at F, and the dotted line $A^{\prime}FD^{\prime}$ is an example. This implies that there may be other $S(\theta)$ functions, i.e., other likelihood function representing different alternative hypothesis, with the same slope at $\theta_{0}$, giving rise to the same RS test statistic. In the literature, this property is known as invariance property of the RS test principle [see Godfrey, (1988, p.70)]. Finally, for both the RS and W tests, the variances can be calculated in a number of ways which are asymptotically equivalent. This can lead to different versions of the test statistics. It is not clear which versions will give better results in finite samples.
@@
RS test was well ahead of its time. It went unnoticed for very many years. It is fair to say that econometricians can claim major credit in recognizing its importance and applying the RS test in several useful contexts and coming up with closed form, neat test statistics. Rao himself acknowledged this fact by writing [see Rao, (2005, p.15)] “I am gratified to see the large number of papers contributed by econometricians on the application of the score statistic to problems in econometrics and the extensions and improvements they have made.” More recently, statisticians are catching up with innovative applications. To obtain a quantitative perception of the influence of Rao, (1948), we plot the yearly citations for the last 75 years in Figure 2. The corresponding cumulative citations are depicted in Figure 3. First thing to note is that the total number of citations in the last 75 years is only 980, apparently a very low number for such a seminal paper. Of course, we need to take into consideration of the fact that there are many papers, especially in the Statistics literature, that use the score test without making any reference to Rao, (1948). Second, there are only a handful citations during the first thirty years, i.e., until around 1978. That was the time econometricians recognized the usefulness of the Rao test principle, and used it in developing several model specification tests. There was another surge in its use after another 30 years, i.e., around 2008, in both the Statistics and Econometrics literature. Finally, from both Figures 2 and 3, it is clear that overall, the number of citations is still going up at an increasing rate, indicating continuing influence of Rao, (1948), as far as the citation numbers go.
@@
$E\left[\sqrt{b_{1}}\right]=0,\quad Var\left[\sqrt{nb_{1}}\right]=6,$ (40)
@@
$P_{N}^{\prime}(\theta_{0})\geq\mu\,P_{N}(\theta_{0}),$ (48)
@@
$P_{n}^{\prime}(\theta_{0})\geq A(N)\,P_{n}(\theta_{0}),$ (49)
@@
Sequential testing procedures that perform interim analyses during the evolution of the experiment, with the goal of obtaining the result earlier than the termination time suggested by the fixed sample analysis due to time or monetary cost considerations or ethical reasons, are easier validated with the use of score-based test statistics rather than the analogues of LR statistics. We may refer to chapters 9-11 of Sen, (1981) for the role of score processes in sequential nonparametrics, where it is mentioned (p.339) “it is comparatively simpler to verify these regularity conditions [i.e., for the score] than those for the likelihood function.”
@@
To diagnose the potential ability of candidate judges and to decide on the selection of a taste panel, each candidate judge is required to perform a prespecified number of sample comparisons. The number of sample comparisons that should be performed by each candidate judge before reaching to a decision on who to include in the taste panel is always a concern.
@@
Bilias, (2000) offered an application of a repeated significance test in a retrospective analysis of the Pennsylvania ‘Reemployment Bonus’ controlled experiments conducted by the US Department of Labor. Their main purpose was to determine whether the offer of a bonus amount to the unemployment insurance (UI) claimants, provided that they find a job with some required permanence within a given period of time, can act as an incentive for more intensive job-seeking with subsequent reduction of the unemployment spells. The response of primary interest is the length of insured unemployment spell and it is assumed that it follows a proportional hazard regression model. The statistic for measuring the effect of the various bonus packages on the duration of insured unemployment relative to the existing scheme is the partial likelihood score statistic. In carrying out sequential analysis, the score statistic is evaluated repeatedly, at different points in chronological time, each time with the available data. The retrospective sequential analysis concluded that the experiment could be concluded earlier than the fixed sample analysis with gains in time and monetary savings.
@@
While discussing the problem in statistical hypothesis testing, Haavelmo, (1944, pp.65-66) stated, “Whatever be the principles by which we choose a “best” critical region of size $\alpha$, the essential thing is that a test is always developed with respect to a given fixed set of possible alternatives $\Omega^{0}$.” Haavelmo called $\Omega^{0}$, the a priori admissible hypotheses and according to him, a test is not robust if we shift our attention to another admissible set $\Omega^{\prime}$ (that may be obtained by extending $\Omega^{0}$ to include new/different alternatives), for which the proposed test has poor size and power properties.
Very often it is difficult to interpret the results of a test applied to a misspecified model. For instance, while testing the significance of some of the regression coefficients in the linear regression models, the results are not easily interpretable when a nonlinear model is the appropriate one [see, White, (1980), Bera and Byron, (1983) and Byron and Bera, (1983)]. In the Statistics and Econometrics literature, most emphasis has been put on the minimization of type-I and type-II error probabilities. There are, however, only a few works that seriously consider the consequences and suggest remedies of misspecifying the a priori admissible hypothesis – which can be called the type-III error.
Note that the model under our a priori admissible hypothesis could be misspecified in a variety of ways. Here we consider only two kinds: distributional and parametric. In the former case, the assumed probability density function differs from the true data generating process (DGP). Kent, (1982) and White, (1982) analyzed this case and suggested a modified version of the RS test that involves adjustment of the variance of the score function. In the parametric misspecification case, the dimension of the assumed parameter space does not match with the true one. Bera and Yoon, (1993) developed a modified RS test that is valid under the local parametric misspecification.
@@
Due to this divergence between $J$ and $K$, and noting that we defined the information matrix $\mathcal{I}(\theta)$ in (10) by taking expectation under $f(y;\theta)$ instead of under the DGP $g(y)$, the standard RS test in (15) is not valid. Let us define an estimator of $\theta$ by maximizing a likelihood function based on the misspecified density $f(y;\theta)$ in place of the unknown DGP $g(y)$. Such an estimator is called quasi-MLE (QMLE). An early reference to QMLE can be found in Koopmans et al., (1950, p.135) [for more on this see Bera et al., (2020)]. We will denote QMLE of $\theta$ (under $H_{0}$) by $\tilde{\theta}$. Kent, (1982) and White, (1982) suggested the following robust form of the RS test statistic for testing the $H_{0}:h(\theta)=c$:
@@
where $H(\theta)=\partial h(\theta)/\partial\theta$, $B(\theta)=K(\theta)^{-1}J(\theta)K(\theta)^{-1}$ and the notation $RS^{*}(D)$ is used to signify robust RS test statistic under distributional misspecification.
Under $H_{0}:h(\theta)=c$, $RS^{*}(D)$ is asymptotcally distributed as $\chi^{2}_{r}$ even under distributional misspecification, that is, when the assumed density $f(y;\theta)$ does not coincide with the true DGP $g(y)$. This approach of finding the asymptotically correct formula for variance has its origin in Koopmans et al., (1950, pp.148-150); for more on this see Bera et al., (2021). Expression (56) can be simplified if the parameter vector $\theta$ $(p\times 1)$ can be partitioned as $\theta=(\gamma^{\prime},\psi^{\prime})^{\prime}$ where $\gamma$ and $\psi$ have dimensions $m$ and $r$, respectively, $m+r=p$, and we test $H_{0}:\psi=\psi_{*}$ (say). Let us also partition the score function $S(\theta)$ and $J(\theta)$ [similarly $K(\theta)$] as